# Towards Anomaly-Aware Pre-Training and Fine-Tuning for Graph Anomaly Detection

### v2

@@ Abstract
Graph anomaly detection (GAD) has garnered increasing attention in recent years, yet remains challenging due to two key factors: (1) label scarcity stemming from the high cost of annotations and (2) homophily disparity at node and class levels. In this paper, we introduce Anomaly-Aware Pre-Training and Fine-Tuning (APF), a targeted and effective framework to mitigate the above challenges in GAD. In the pre-training stage, APF incorporates node-specific subgraphs selected via the Rayleigh Quotient, a label-free anomaly metric, into the learning objective to enhance anomaly awareness. It further introduces two learnable spectral polynomial filters to jointly learn dual representations that capture both general semantics and subtle anomaly cues. During fine-tuning, a gated fusion mechanism adaptively integrates pre-trained representations across nodes and dimensions, while an anomaly-aware regularization loss encourages abnormal nodes to preserve more anomaly-relevant information. Furthermore, we theoretically show that APF tends to achieve linear separability under mild conditions. Comprehensive experiments on 10 benchmark datasets validate the superior performance of APF in comparison to state-of-the-art baselines.
@@ 1 Introduction
Graph anomaly detection (GAD) aims to identify a small but significant portion of instances, such as abnormal nodes, that deviate significantly from the standard, normal, or prevalent patterns within graph-structured data GADSurvey . The detection of these anomalies is crucial for various scenarios, such as financial fraud in transaction networks FraudSurvey1 ; FraudSurvey2 , fake news and bots in social media FakeNewsSurvey ; BotSurvey , and sensor faults in IoT networks IoTSurvey1 ; IoTSurvey2 . Given their strong ability to model relational structure, graph neural networks (GNNs) have recently emerged as a leading choice for tackling GAD.
@@
Despite notable progress, most existing GAD methods CARE-GNN ; PCGNN ; BWGNN ; GHRN are not tailored for label-scarce scenarios, leading to suboptimal performance in real-world deployments where annotations are costly. Recent semi-supervised attempts, such as pseudo-labeling UPS ; ConsisGAD and synthetic sample generation CGenGA ; GGAD , seek to mitigate this but often suffer from instability due to the inherent uncertainty and confirmation bias that compromise performance UPS . In contrast, the pre-training-and-fine-tuning paradigm has shown great promise in label-scarce learning across CV SimCLRv2 ; SSLEL , NLP BERT ; TAPT , and graph learning GRACE ; DEGL . A recent study GPT further reveals that general-purpose graph pre-training strategies DGI ; GraphMAE can match or even outperform GAD-specific methods under limited supervision. Despite their strong potential, existing graph pre-training-and-fine-tuning frameworks are primarily designed to extract task-agnostic semantic knowledge. As such, they fail to address GAD’s unique challenges, falling short in capturing anomaly-relevant cues and leaving their adaptation to GAD an unsolved yet pressing task.
@@
In the context of GAD, global homophily, the same-class edge ratio over the entire graph, tends to decrease due to challenges like camouflage CARE-GNN , making it an intuitive yet common anomaly indicator. Nonetheless, we highlight that local homophily, the class consistency within each node’s neighborhood, reveals more nuanced disparity patterns, capturing localized yet often overlooked anomaly signals. As illustrated in Figures 1(a) and 1(b), these disparities manifest at two granularities: (1) node-level disparity represents the high variations in local homophily across individual nodes and (2) class-level disparity refers to lower local homophily for abnormal nodes. Most existing approaches are built around global homophily and employ uniform processing schemes, such as edge reweighting H2-FDetector ; GHRN or spectral filtering BWGNN ; AMNet . However, such globally uniform designs lack node-adaptive mechanisms necessary to accommodate the structural diversity of individual nodes, resulting in inconsistent anomaly distinguishability across nodes in different local homophily groups, as further evidenced in Figures 1(c) and 1(d). These limitations jointly highlight a critical challenge: How can we devise a GAD-specific pre-training-and-fine-tuning framework that effectively handles dual-granularity local homophily disparity?
@@
To enable granularity-adaptive fine-tuning, APF employs a gated fusion network that adaptively combines pre-trained representations at both the node and dimension levels. An anomaly-aware regularization loss is further introduced to encourage abnormal nodes to retain more anomaly-relevant information from pre-trained representations than normal nodes. Together, these components explicitly address local homophily disparity, by aligning the fine-tuning with the homophily disparity at node and class levels under label-guided optimization. Theoretical analysis further shows that, under mild conditions, APF tends to achieve linear separability across all nodes.
@@ 2 Preliminary
In this section, we briefly introduce notations and key concepts. Detailed preliminaries and related works are provided in Appendix A and Appendix B, respectively.
@@ 3.1 Anomaly-Aware Pre-Training
Most existing graph pre-training strategies DGI ; CCA-SSG ; BGRL ; GraphMAE ; BLNN are designed to optimize task-agnostic objectives. As a result, the learned representations primarily encode general semantic knowledge. In contrast, the goal of GAD is to capture subtle anomaly cues that differentiate minority abnormal instances. However, many of these cues, such as relation camouflage CARE-GNN ; PCGNN , are inherently label-dependent, making them difficult to capture under the label-free pre-training.
This motivates the incorporation of effective label-free anomaly measures into the learning objective to guide the representation toward anomaly-aware semantics. Recent findings reveal that the existence of anomalies induces the ‘right-shift’ phenomenon BWGNN ; GHRN , where spectral energy distribution concentrates more on high frequencies. The corresponding accumulated spectral energy can be quantified by the Rayleigh Quotient RQ :
@@
where $\boldsymbol{x}\in\mathbb{R}^{n}$ represents a general graph signal and $\hat{\boldsymbol{x}}=\boldsymbol{U}^{T}\boldsymbol{x}$ denotes its projection in spectral space. By definition, Rayleigh Quotient also quantifies global structural diversity, yielding higher values in the presence of relation camouflage. The following lemma further illustrates the relationship between the Rayleigh Quotient and anomaly information.
@@
The Rayleigh Quotient $RQ(x,L)$, which represents the accumulated spectral energy of a graph signal, increases monotonically with the anomaly degree. BWGNN
Given this, Rayleigh Quotient stands out as a promising label-free anomaly metric. To guide each node $v_{i}$ in capturing its potential anomaly cues, we employ MRQSampler UniGAD to extract its 2-hop sub-tree $\mathcal{G}_{i}^{RQ}$. Each sub-tree is selected to maximize the Rayleigh Quotient, thereby preserving as much anomaly-relevant signal and structural diversity as possible.
@@
Beyond the learning objective, the architectural modules for encoding anomaly-relevant information are equally critical. Recent studies highlight that capturing high-frequency components is essential for modeling heterophilic patterns FAGCN ; ACM ; EvenNet . Motivated by this, we complement the conventional low-pass encoder with an additional high-pass encoder to better capture anomaly cues. To facilitate flexible spectral encoding, we adopt the learnable $K$-order Chebyshev polynomial ChebNet2 and restrict it to fit only low-pass and high-pass filters, denoted by $g_{L}(\cdot)$ and $g_{H}(\cdot)$, following prior work PolyGCL :
$g_{L}(\hat{\boldsymbol{L}})=\sum_{k=0}^{K}w_{k}^{L}T_{k}(\hat{\boldsymbol{L}}),\,g_{H}(\hat{\boldsymbol{L}})=\sum_{k=0}^{K}w_{k}^{H}T_{k}(\hat{\boldsymbol{L}}),$ (4)
@@
where $\boldsymbol{\gamma}=(\gamma_{0},\cdots,\gamma_{M})$ is the shared learnable parameter with $\gamma_{0}=\gamma_{0}^{L}=\gamma_{0}^{H}$. As per PolyGCL , we have $\gamma_{i}^{H}\leq\gamma_{i+1}^{H}$ and $\gamma_{i}^{L}\geq\gamma_{i+1}^{L}$, thus guarantee the high-pass/low-pass property for $g_{L}(\cdot)$/$g_{H}(\cdot)$. Given such, our low-pass and high-pass encoders are formulated as:
@@
where $\boldsymbol{Z}_{L},\boldsymbol{Z}_{H}\in\mathbb{R}^{n\times e}$ denote the low-pass and high-pass node representations respectively. $f_{\theta_{L}}(\cdot),f_{\theta_{H}}(\cdot)$ represent the learnable MLP for each filter.
@@
We build our anomaly-aware pre-training framework upon DGI DGI , owing to its efficiency and proven ability to capture semantic structure. To encode general semantic knowledge, we follow the standard DGI formulation using low-pass encoders. To capture subtle anomaly cues, we further extend this paradigm by maximizing the mutual information between each node and the summary of its Rayleigh Quotient-guided sub-tree, computed from high-pass encoders.
Let $\tilde{\boldsymbol{Z}}_{L}$ and $\tilde{\boldsymbol{Z}}_{H}$ denote the negative samples generated from randomly shuffled inputs DGI . Our dual-objective pre-training loss is formulated as:
@@
where $\boldsymbol{s}^{L}=\frac{1}{n}\sum_{i=1}^{n}\boldsymbol{Z}_{i}^{L}$ is the global summary and $\boldsymbol{s}_{i}^{H}=\frac{1}{|\mathcal{G}_{i}^{RQ}|}\sum_{v_{j}\in\mathcal{G}_{i}^{RQ}}\boldsymbol{Z}_{j}^{H}$ is the summary over the Rayleigh Quotient-based sub-tree of node $v_{i}$. The discriminator is defined as $\mathcal{D}(\boldsymbol{z},\boldsymbol{s})=\sigma(\boldsymbol{z}^{\top}\boldsymbol{W}\boldsymbol{s})$. This dual-objective formulation enables the encoders to jointly capture task-agnostic semantic knowledge and node-specific structural disparities, yielding more informative representations for downstream anomaly detection.
@@ 3.2 Granularity-Adaptive Fine-Tuning
After pre-training, we aim to develop a node-adaptive fusion mechanism that selectively combines task-agnostic semantic knowledge ($\boldsymbol{Z}_{L}$) and node-specific structural disparities ($\boldsymbol{Z}_{H}$) from the frozen pre-trained representations. This fusion is designed to better respond to local homophily variations across nodes under label-guided learning. Beyond trivial node-wise fusion, prior studies JacobiConv ; AdaGNN ; GFS highlight that different feature dimensions contribute unequally to downstream tasks, motivating a dimension-aware fusion design. To this end, we introduce a coefficient matrix $\boldsymbol{C}\in[0,1]^{n\times e}$ to combine $\boldsymbol{Z}_{L}$ and $\boldsymbol{Z}_{H}$ at node and dimension levels:
@@
where $\odot$ denotes the Hadamard product, and $\boldsymbol{Z}$ is the resulting fused representation passed to the classifier. A naive approach to learn $\boldsymbol{C}$ is treating it as free parameters JacobiConv , but this leads to excessive overhead ($\mathcal{O}(n\times e)$), inefficient learning under sparse supervision, and neglect of input semantics. To alleviate this, we introduce a lightweight Gated Fusion Network (GFN) that generates coefficients based on the input features:
@@
where $\boldsymbol{W}_{c}\in\mathbb{R}^{d\times e}$ is a learnable matrix and $\boldsymbol{b}_{c}\in\mathbb{R}^{e}$ is a bias term.
The advantages of GFN over direct optimization are multifold: (1) GFN reduces the learnable parameter complexity of $\boldsymbol{C}$ from $\mathcal{O}(n\times e)$ to $\mathcal{O}((d+1)\times e)$, where $d,e\ll n$; (2) GFN allows sparse supervision to update the entire coefficient matrix, while direct optimization only affects labeled rows; (3) GFN leverages raw input features, which encode valuable anomaly-relevant attributes GADBench .
@@
Given the fused representations $\boldsymbol{Z}$ from Eq. (9), we employ a two-layer MLP to predict the label $\hat{y}_{i}$ for each labeled node $v_{i}$. The model is optimized using the standard binary cross-entropy loss:

### v3

@@ Abstract
Graph anomaly detection (GAD) has garnered increasing attention in recent years, yet remains challenging due to two key factors: (1) label scarcity stemming from the high cost of annotations and (2) homophily disparity at node and class levels. In this paper, we introduce Anomaly-Aware Pre-Training and Fine-Tuning (APF), a targeted and effective framework to mitigate the above challenges in GAD. In the pre-training stage, APF incorporates node-specific subgraphs selected via the Rayleigh Quotient, a label-free anomaly metric, into the learning objective to enhance anomaly awareness. It further introduces two learnable spectral polynomial filters to jointly learn dual representations that capture both general semantics and subtle anomaly cues. During fine-tuning, a gated fusion mechanism adaptively integrates pre-trained representations across nodes and dimensions, while an anomaly-aware regularization loss encourages abnormal nodes to preserve more anomaly-relevant information. Furthermore, we theoretically show that APF tends to achieve linear separability under mild conditions. Comprehensive experiments on 10 benchmark datasets validate the superior performance of APF in comparison to state-of-the-art baselines. The code is available at https://github.com/Cloudy1225/APF.
@@ 1 Introduction
Graph anomaly detection (GAD) aims to identify a small but significant portion of instances, such as abnormal nodes, that deviate significantly from the standard, normal, or prevalent patterns within graph-structured data (Qiao et al., 2025b). The detection of these anomalies is crucial for various scenarios, such as financial fraud in transaction networks (Cheng et al., 2025), fake news in social media (Aïmeur et al., 2023), and sensor faults in IoT networks (Gaddam et al., 2020). Given their strong ability to model relational structure, graph neural networks (GNNs) have recently emerged as a leading choice for tackling GAD.
@@
Despite notable progress, most existing GAD methods (Dou et al., 2020; Tang et al., 2022; Zhuo et al., 2024; Zheng et al., 2025a) are not tailored for label-scarce scenarios, leading to suboptimal performance in real-world deployments where annotations are costly. Recent semi-supervised attempts, such as pseudo-labeling (Rizve et al., 2021; Chen et al., 2024b) and synthetic sample generation (Ma et al., 2024; Qiao et al., 2024), seek to mitigate this but often suffer from instability due to the inherent uncertainty and confirmation bias that compromise performance (Rizve et al., 2021). In contrast, the pre-training and fine-tuning paradigm has shown great promise in label-scarce learning across CV (Chen et al., 2020; Nandam et al., 2025), NLP (Devlin et al., 2019; Shi et al., 2023), and graph learning (Zhu et al., 2020; Ju et al., 2024). A recent study (Cheng et al., 2024) further reveals that general-purpose graph pre-training strategies (Veličković et al., 2019; Hou et al., 2022) can match or even outperform GAD-specific methods under limited supervision. Despite their strong potential, existing graph pre-training frameworks are primarily designed to extract task-agnostic semantic knowledge. As such, they fail to address GAD’s unique challenges, falling short in capturing anomaly-relevant cues and leaving their adaptation to GAD an unsolved yet pressing task.
@@
In GAD, global homophily, the intra-class edge ratio over the entire graph, tends to decrease due to challenges like camouflage (Dou et al., 2020), making it an intuitive yet common anomaly indicator. Nonetheless, we highlight that local homophily, the class consistency within each node’s neighborhood, reveals more nuanced disparity patterns, capturing localized yet often overlooked anomaly signals. As illustrated in Figures 1(a) and 1(b), these disparities manifest at two granularities: (1) node-level disparity represents the high variations in local homophily across individual nodes and (2) class-level disparity refers to lower local homophily for abnormal nodes. Most existing approaches are built around global homophily and employ uniform processing schemes, such as edge reweighting (Shi et al., 2022; Gao et al., 2023) or spectral filtering (Tang et al., 2022; Zheng et al., 2025a). However, such globally uniform designs lack node-adaptive mechanisms necessary to accommodate the structural diversity of individual nodes, resulting in inconsistent anomaly distinguishability across nodes in different local homophily groups, as further evidenced in Figures 1(c) and 1(d). These limitations jointly highlight a critical challenge: How can we devise a GAD-specific pre-training and fine-tuning framework that effectively mitigates dual-granularity local homophily disparity?
@@
To enable granularity-adaptive fine-tuning, APF employs a gated fusion network that adaptively combines pre-trained representations at both the node and dimension levels. An anomaly-aware regularization loss is further introduced to encourage abnormal nodes to retain more anomaly-relevant information from pre-trained representations than normal nodes. Together, these components explicitly address local homophily disparity, by aligning the fine-tuning with the homophily disparity at node and class levels under label-guided optimization. Theoretical analysis further shows that APF tends to achieve linear separability across all nodes.
@@ 2 Preliminary
In this section, we briefly introduce notations and key concepts. Detailed preliminaries and related works are provided in Appendix B and Appendix C, respectively.
@@ 3.1 Anomaly-Aware Pre-Training
Most existing graph pre-training strategies (Veličković et al., 2019; Hou et al., 2022; Liu et al., 2024d) are designed to optimize task-agnostic objectives. As a result, the learned representations primarily encode general semantic knowledge. In contrast, the goal of GAD is to capture subtle anomaly cues that differentiate minority abnormal instances. However, many of these cues, such as relation camouflage (Dou et al., 2020), are inherently label-dependent, making them difficult to capture under the label-free pre-training.
This motivates the incorporation of effective label-free anomaly measures into the learning objective to guide the representation toward anomaly-aware semantics. Recent findings reveal that the existence of anomalies induces the ‘right-shift’ phenomenon (Tang et al., 2022; Gao et al., 2023; Dong et al., 2024), where spectral energy distribution concentrates more on high frequencies. The corresponding accumulated spectral energy can be quantified by the Rayleigh Quotient (Horn and Johnson, 2012):
@@
where $\boldsymbol{x}\in\mathbb{R}^{n}$ represents a general graph signal and $\hat{\boldsymbol{x}}=\boldsymbol{U}^{T}\boldsymbol{x}$ denotes its projection in spectral space. Crucially, the Rayleigh Quotient measures the inconsistency between node attributes and the local graph structure (i.e., graph smoothness). A high value indicates that connected nodes have dissimilar features, which allows RQ to capture both attribute anomalies (where features deviate from neighbors) and structural anomalies (e.g., camouflaged edges connecting dissimilar nodes) (Tang et al., 2022). We provide a detailed discussion on how RQ captures these different anomaly types in Appendix K. The following lemma further illustrates the relationship between the Rayleigh Quotient and anomaly information.
@@
The Rayleigh Quotient $RQ(x,L)$, which represents the accumulated spectral energy of a graph signal, increases monotonically with the anomaly degree. (Tang et al., 2022)
Given this, Rayleigh Quotient stands out as a promising label-free anomaly metric. To guide each node $v_{i}$ in capturing its potential anomaly cues, we employ MRQSampler (Lin et al., 2024) to extract its 2-hop subgraph $\mathcal{G}_{i}^{RQ}$. Each subgraph is selected to maximize the Rayleigh Quotient, thereby preserving as much structural diversity and anomaly-relevant signal as possible (Lin et al., 2024).
@@
Beyond the learning objective, the architectural modules for encoding anomaly-relevant information are equally critical. Recent studies highlight that capturing high-frequency components is essential for modeling heterophilic patterns (Bo et al., 2021; Luan et al., 2022; Lei et al., 2022). Motivated by this, we complement the conventional low-pass encoder with an additional high-pass encoder. This dual-branch design allows us to simultaneously capture general semantic patterns (via low-pass filters) and subtle anomaly cues (via high-pass filters), the latter of which are often smoothed out by standard GNNs. We provide a detailed discussion on the motivation and spectral guarantees of this design in Appendix L. To facilitate flexible spectral encoding, we adopt the learnable $K$-order Chebyshev polynomial (He et al., 2022) and restrict it to fit only low-pass and high-pass filters, denoted by $g_{L}(\cdot)$ and $g_{H}(\cdot)$, following prior work Chen et al. (2024a):
[…]
@@
where $\boldsymbol{\gamma}=(\gamma_{0},\cdots,\gamma_{M})$ is the shared learnable parameter with $\gamma_{0}=\gamma_{0}^{L}=\gamma_{0}^{H}$. As per Chen et al. (2024a), we have $\gamma_{i}^{H}\leq\gamma_{i+1}^{H}$ and $\gamma_{i}^{L}\geq\gamma_{i+1}^{L}$, thus guarantee the high-pass/low-pass property for $g_{L}(\cdot)$$g_{H}(\cdot)$. Given such, our low-pass and high-pass encoders are formulated as:
@@
where $\boldsymbol{Z}_{L},\boldsymbol{Z}_{H}\in\mathbb{R}^{n\times e}$ denote the low-pass and high-pass node representations respectively. $f_{\theta_{L}}(\cdot),f_{\theta_{H}}(\cdot)$ represent the learnable MLP for each filter. To ensure consistent scaling for subsequent fusion, both $\boldsymbol{Z}_{L}$ and $\boldsymbol{Z}_{H}$ are standardized with zero mean and unit variance.
@@
Unlike conventional objectives that only preserve task-agnostic semantics, our optimization explicitly integrates anomaly awareness. We build the pre-training framework upon DGI (Veličković et al., 2019) owing to its efficiency and effectiveness: we retain its standard low-pass branch to encode generic semantic structure, and further incorporate an anomaly-sensitive objective that maximizes the mutual information between each node and the summary of its Rayleigh Quotient-guided subgraph, computed by high-pass encoders. Let $\tilde{\boldsymbol{Z}}_{L}$ and $\tilde{\boldsymbol{Z}}_{H}$ denote negative samples generated from randomly shuffled inputs (Veličković et al., 2019). Our pre-training loss is:
@@
where $\boldsymbol{s}^{L}=\frac{1}{n}\sum_{i=1}^{n}\boldsymbol{Z}_{i}^{L}$ is the global semantic summary and $\boldsymbol{s}_{i}^{H}=\frac{1}{|\mathcal{G}_{i}^{RQ}|}\sum_{v_{j}\in\mathcal{G}_{i}^{RQ}}\boldsymbol{Z}_{j}^{H}$ is the anomaly-aware summary over the Rayleigh Quotient-based subgraph of node $v_{i}$. The discriminator is defined as $\mathcal{D}(\boldsymbol{z},\boldsymbol{s})=\sigma(\boldsymbol{z}^{\top}\boldsymbol{W}\boldsymbol{s})$. By jointly optimizing these two complementary objectives, our framework departs from existing pre-training methods and introduces a principled way to capture task-agnostic semantic knowledge and node-specific anomaly cues, yielding more informative representations for downstream anomaly detection.
@@ 3.2 Granularity-Adaptive Fine-Tuning
After pre-training, we aim to develop a node-adaptive fusion mechanism that selectively combines task-agnostic semantic knowledge ($\boldsymbol{Z}_{L}$) and node-specific structural disparities ($\boldsymbol{Z}_{H}$) from the frozen pre-trained representations. This fusion is designed to better respond to local homophily variations across nodes under label-guided learning. Beyond trivial node-wise fusion, prior studies (Wang and Zhang, 2022; Dong et al., 2021; Zheng et al., 2025b) highlight that different feature dimensions contribute unequally to downstream tasks, motivating a dimension-aware fusion design. To this end, we introduce a coefficient matrix $\boldsymbol{C}\in[0,1]^{n\times e}$ to combine $\boldsymbol{Z}_{L}$ and $\boldsymbol{Z}_{H}$ at node and dimension levels:
@@
where $\odot$ denotes the Hadamard product, and $\boldsymbol{Z}$ is the resulting fused representation passed to the classifier. A naive approach to learn $\boldsymbol{C}$ is treating it as free parameters (Wang and Zhang, 2022), but this leads to excessive overhead ($\mathcal{O}(n\times e)$), inefficient learning under sparse supervision, and neglect of input semantics. To alleviate this, we introduce a lightweight Gated Fusion Network (GFN) that generates coefficients based on the input features:
@@
where $\boldsymbol{W}_{c}\in\mathbb{R}^{d\times e}$ is a learnable matrix and $\boldsymbol{b}_{c}\in\mathbb{R}^{e}$ is a bias term, $\sigma(\cdot)$ is sigmoid function. The advantages of GFN over direct optimization are multifold: (1) GFN reduces the learnable parameter complexity of $\boldsymbol{C}$ from $\mathcal{O}(n\times e)$ to $\mathcal{O}((d+1)\times e)$, where $d,e\ll n$; (2) GFN allows sparse supervision to update the entire coefficient matrix, while direct optimization only affects labeled rows; (3) GFN leverages raw input features, which encode valuable anomaly-relevant attributes (Tang et al., 2023).
@@
Given the fused representations $\boldsymbol{Z}$ from Eq. 9, we employ a two-layer MLP to predict the label $\hat{y}_{i}$ for each labeled node $v_{i}$. The model is optimized using the standard binary cross-entropy loss: